### v3

Abstract
Using a theorem of Jackiw and Pi expressing the delicate balance of the spin and the orbital momentum, we systematically classify the flat-space massless Lagrangian quantum field theories that are invariant under the global conformal group $SO(\texttt{D},2)$. We recover in a uniform way the facts that scalars and spinors are invariant in any dimension, and that gauge p-tensors are invariant only in $2\texttt{p}+2$ dimensions. This case includes the Maxwell theory in 4 dimensions and the Kalb-Ramond 2-forms theory in 6 dimensions. We then construct two new classes of Lagrangians extending the Avdeev-Chizhov self-dual tensor model to higher dimensions, one class using a symmetric metric and the other a skew metric in internal space. Finally, we prove in the same uniform way that both classes are conformal invariant in any even dimension. In 4 dimensions, these self-dual tensors naturally couple to the chiral Fermions of the standard model.
keywords:
Conformal invariance, Antisymmetric tensor fields , Avdeev-Chizhov , QFT
journal: J Phys A
[…]
1 Introduction
The aim of this article is to study the conformal invariance of Lagrangians describing antisymmetric tensor fields in any even dimension using the elegant theorem formulated by Jackiw and Pi (2011).
Despite the fact that the conformal group plays a crucial role in the superstring models which cover all dimensions from 2 to 10, there is, to our knowledge, no systematic general classification of the flat space conformally invariant Lagrangian field theories. To fill this gap we provide here a direct and uniform proof for the known cases of the scalars, spinors vectors, and the Kalb-Ramond gauge tensors (KR) theories and then extend our proof to two new generalizations in D dimensions of the Avdeev and Chizhov (1994b, a) self-dual tensors. The first class (AC) uses a symmetric metric in internal space and the second class (CP), an anti-symmetric metric. We call it (CP) because such Lagrangians flip sign under (C) charge or (P) parity flips, but are invariant under Landau (CP) symmetry. We prove that the massless scalars and spinors are conformal invariant in any dimension, that the (KR) Lagrangians describing antisymmetric tensors with p indices are conformal invariant only in dimension $\texttt{D}=2\texttt{p}+2$ and then the new result that the (AC) and (CP) Lagrangians describing self-dual tensors with p indices are conformal invariant in any even dimension $\texttt{D}=2\texttt{p}$. Notice that Witten (2004) and followers study an entirely different kind of theories where rather than the tensor $T$ itself, its curvature $dT$ is self-dual. Such models are not part of our classification because they do not admit a Lagrangian. On the other hand, these authors do not consider our new self dual classes (AC) or (CP).
To understand the literature dealing with conformal invariance, three symmetry groups must be carefully distinguished: (A) the 1 dimensional global scale transformations, (B) the $(\texttt{D}+2)(\texttt{D}+1)/2$ dimensional global conformal group $SO(\texttt{D},2)$ which combines the dilatations, the $\texttt{D}(\texttt{D}-1)/2$ Lorentz rotations, the $D$ Poincaré translations, and the $D$ boosts, in total 15 parameters in $\texttt{D}=4$ dimensions, and (C) the infinite dimensional Weyl local scale transformations which apply to curved space.
Although they are undoubtedly different, these three groups are strongly related. By definition, (B) and (C) contain (A). Furthermore the three groups have interconnected applications in physics. For example, one objective of the seminal physics reports by Nakayama (2015) was to establish that global scale invariance (A), combined with unitarity, implies conformal invariance (B), which, according to Farnsworth et al. (2017), in turn implies Weyl invariance (C). And in the sixties, as recalled in the Nakayama’s introduction, Migdal could not easily convince the great Bogolioubov that conformal invariance (B) was more powerful than scale invariance (A).
However, full equivalence is not established: for example, Karananas and Monin (2016) show that not all scalar conformal invariant theories (B) can be made locally scale invariant (C). Several other studies, for example El-Showk et al. (2011); Wu (2017); Nakayama (2017); Lee and Zheng (2021), show that there exist (A) scale invariant and unitary theories which are not (B) conformally invariant.
A contrario, the interplay between these three groups is sometimes overlooked. For example, in a paper titled ’the smile of the Cheshire cat’, Shapiro (2024) has studied an antisymmetric tensor model in curved space invariant under local scale invariance (C: the cat). In the flat space limit, although local scale invariance disappears, the model maintains wonderland properties and Shapiro compares this phenomenon to the Cheshire cat, because he does not notice that the tail of the cat remains visible. Indeed, in that limit, his model coincides with the Avdeev-Chizhov model (AC). Therefore it is not only scale invariant (A: the smile of the cat) but also invariant under the global conformal group (B: the tail of the cat). Although we have no proof, it is not inconceivable that, as in this model of Shapiro, the flat space limit of any locally scale invariant field theory (C) should be globally conformal invariant (B).
Let us now sketch the organization of our analysis. We only consider Minkowski flat space. In dimension D any Lorentz invariant free field Lagrangian $\mathcal{L}(\Phi,\partial\Phi)$ is automatically scale invariant. We just need to choose the canonical dimension d of each field such that the dimension of each propagator has total dimension $D$, counting as 1 the dimension of the derivatives. In a Bosonic model with two derivatives, like the scalar or the Maxwell theory, $\texttt{D}=2\texttt{d}+2$. In a Fermionic Dirac Lagrangian with one derivative $\texttt{D}=2\texttt{d}+1$. In a Bosonic quartic model (appendix B), $\texttt{D}=2\texttt{d}+4$. Couplings can then be added as long as the coupling constants are dimensionless.
Naïvely, it may then seem that any globally (A) scale invariant field theory should be (B) conformal invariant, but this is not true. As formalized by Jackiw and Pi (2011), the problem lies in the influence of the conformal transformations on field derivatives because the field derivatives are not tensorial, meaning that their transformation rules depend on the derivatives of the coordinate transformations. This affects the apparent orbital momentum and spin of the particles. But for the theory to be invariant, these two modifications must exactly compensate each other, otherwise the apparent rotational energy of the fields would not be conserved. This constrains the structure of the field propagators and of all couplings involving field derivatives as they must exquisitely match the internal spin of the particles.
In section 2, we recall the theorem of Jackiw and Pi (2011) and systematically prove that the classic scalar and spin-half free Lagrangians are conformal invariant in any dimension, that the Maxwell Lagrangian is invariant only in $\texttt{D}=4$ dimensions (Bateman (1910)) and that covariant derivatives and Yang-Mills-Higgs couplings are conformal invariant in $\texttt{D}=4$ because the coupling constants are dimensionless. We also discuss the spin 3/2 supergravity field.
In section 3, we prove that the free scale invariant (A) and gauge invariant massless Kalb-Ramond (KR) p-form Lagrangians $\mathcal{L}=({}^{*}dT)\wedge(dT)$ are conformal invariant (B) only in dimension $\texttt{D}=2\texttt{p}$ (see below for the notations). Although we have here provided the benefit of a systematic proof, the results thus far are known. We then turn to new results. We study the Avdeev and Chizhov (1994b, a) Lagrangian (AC), which describes self-dual $2$-forms in $4$ dimensions, generalize it to D dimensions, and further generalize it to a new anti-symmetric (CP) model, first encountered in our recent study of the $SU(2/1)$ superalgebraic structure of the electroweak interactions (Thierry-Mieg and Jarvis (2021)). We prove that the (AC) and the (CP) models are both conformally invariant (B) in any even dimension.
In the short section 4, we discuss the physical content of antisymmetric tensor field theories and count the degrees of freedom in D dimensions. We also show that in $\texttt{D}=4$ dimensions, the triple scalar-vector-tensor coupling $\Phi F^{\mu\nu}T_{\mu\nu}$ which involves derivatives ($F_{\mu\nu}=\partial{{}_{[}\mu}A_{\nu]}$) is conformal invariant.
In the appendices, we analyze 2 additional toy models discussed in the literature. We show in a simpler way that the curious gauge fixed Maxwell model of El-Showk et al. (2011) is conformal invariant, but that contrary to section C.3 of Lee and Zheng (2021), the scale invariant quartic Maxwell model $F^{\mu\nu}\Box F_{\mu\nu}$ is not conformal invariant.
2 Conformal invariance
A basis of the generators of the $SO(D,2)$ conformal Lie algebra in D dimensions with Minkowski metric $g_{\mu\nu}$ is provided by the $SO(D-1,1)$ Lorentz transformations $M_{\mu\nu}$, the dilatation $\Delta$, the translations $P_{\mu}$ and the conformal boosts $K_{\mu}$ with commutation relations (see for example El-Showk et al. (2011), equation A.1).
$\begin{array}[]{c}[M_{\mu\nu},M_{\rho\sigma}]=ig_{\nu\rho}M_{\mu\sigma}-ig_{\nu\sigma}M_{\mu\rho}-ig_{\mu\rho}M_{\nu\sigma}+ig_{\mu\sigma}M_{\nu\rho}\;,\\ \;[M_{\mu\nu},P_{\rho}]=ig_{\nu\rho}P_{\mu}-ig_{\mu\rho}P_{\nu}\;,\;\;[M_{\mu\nu},K_{\rho}]=ig_{\nu\rho}K_{\mu}-ig_{\mu\rho}K_{\nu}\;,\\ \;[P_{\mu},K_{\nu}]=+2i(g_{\mu\nu}\Delta-M_{\mu\nu})\;,\\ \;[\Delta,M_{\mu\nu}]=0\;,\;\;[\Delta,P_{\mu}]=-iP_{\mu}\;,\;\;]\;,\;\;[\Delta,K_{\mu}]=-iK_{\mu}\;.\end{array}$ (1)
The algebra acts on fields $\Phi$ with canonical scaling dimension d as
$\begin{array}[]{c}\delta\Phi=(\epsilon^{\mu\nu}(-i(x_{\mu}\partial_{\nu}-x_{\nu}\partial_{\mu})+\Sigma_{\mu\nu})-i\;\epsilon^{\mu}\partial_{\mu}\\ -i\;\epsilon^{\prime\mu}(2x_{\mu}\texttt{d}+2ix^{\alpha}\Sigma_{\alpha\mu}+2x_{\mu}x^{\alpha}\partial_{\alpha}-X^{2}\partial_{\mu})-i\;\epsilon(x^{\mu}\partial_{\mu}+\texttt{d}))\Phi\end{array}$ (2)
where $\epsilon^{\mu\nu}$ parameterize the Lorentz transformations $M_{\mu\nu}$  and $\Sigma_{\mu\nu}$ acts on the spin of the field $\Phi$ (see the definitions below), $\epsilon^{\mu}$ parameterize the Poincaré translations $P_{\mu}$, $\epsilon^{\prime\mu}$ the conformal boosts $K_{\mu}$  and finally $\epsilon$ the dilatation $\Delta$. The caveat is that the action of the boosts on the fields is non trivial.
Jackiw and Pi (2011), elaborating on the earlier work of Callan et al. (1970), have proposed a very elegant criterion to assess the conformal invariance of a globally scale invariant and Poincaré invariant Lagrangian. The necessary and sufficient condition is that the field virial
$\begin{array}[]{c}\par V_{\sigma}=\displaystyle{\frac{\partial\mathcal{L}}{\partial(\partial_{\alpha}\Phi)}}\;(\texttt{d}\;g_{\alpha\sigma}-\Sigma_{\alpha\sigma})\;\Phi\end{array}$ (3)
summed over all fields $\Phi$ should be a divergence,
$\begin{array}[]{c}V_{\sigma}=\partial^{\alpha}\omega_{\sigma\alpha}\;.\end{array}$ (4)
In this definition, $\mathcal{L}$ is the Lagrangian, $g_{\mu\nu}$ is the metric, d is the canonical dimension of the fields, $\omega$ is an arbitrary local polynomial in the fields and their derivatives, and $\Sigma$ denotes the Lorentz transformation generator acting on the internal spin of the field $\Phi$.
To use this criterion, we need to compute d and to normalize $\Sigma$ properly. On a scalar field, $\texttt{D}=2\texttt{d}+2$ and $\Sigma$ vanishes. On a chiral spinor, $\texttt{D}=2\texttt{d}+1$ and the rotations are represented by the Dirac matrices. On a tensor, $\texttt{D}=2\texttt{d}+2$ and each vector index rotates like the coordinates.
$\begin{array}[]{c}\Sigma_{\alpha\beta}(\phi)=0\;,\;\;\;\Sigma_{\alpha\beta}(A_{\mu})=-g_{\beta\mu}A_{\alpha}+g_{\alpha\mu}A_{\beta},\\ \Sigma_{\alpha\beta}(\psi)=\frac{1}{4}(\gamma_{\alpha}\gamma_{\beta}-\gamma_{\beta}\gamma_{\alpha})\;\psi\;,\\ \Sigma_{\alpha\beta}(T_{\mu\nu...})=(-g_{\beta\mu}T_{\alpha\nu...}+g_{\alpha\mu}T_{\beta\nu...})+\\ +(g_{\beta\nu}T_{\mu\alpha...}-g_{\alpha\nu}T_{\mu\beta...})+...\end{array}$ (5)
The rotation of a fully antisymmetric p-tensor $T$, i.e. a p-form, involves p pairs of terms. As a first test, we compute the field virial $V_{\sigma}$ for a real free massless scalar field
$\begin{array}[]{c}\mathcal{L}=-\frac{1}{2}\partial_{\mu}\Phi\;\partial^{\mu}\Phi\;,\\ V_{\sigma}(\Phi)=\displaystyle{\frac{\partial\mathcal{L}}{\partial(\partial_{\alpha}\Phi)}}\;(\texttt{d}\;g_{\alpha\sigma})\;\Phi=-\frac{1}{2}\;\texttt{d}\;\partial_{\sigma}\Phi^{2}\;.\end{array}$ (6)
We found a divergence, therefore the scalar Lagrangian is conformal invariant in any dimension. We now try the massless spinors
$\begin{array}[]{c}\mathcal{L}=i{\overline{\psi}}\gamma^{\mu}\partial_{\mu}\psi\;,\\ V_{\sigma}(\psi)=\displaystyle{\frac{\partial\mathcal{L}}{\partial(\partial_{\alpha}\psi_{L})}}\;(\texttt{d}\;g_{\alpha\sigma}\;-\Sigma_{\alpha\sigma})\;\psi\\ =i{\overline{\psi}}\gamma_{\sigma}(\texttt{d}-(\texttt{D}-1)/2)\psi=0\;.\end{array}$ (7)
Here, the canonical dimension d where $\texttt{d}=(\texttt{D}-1)/2$ is exactly compensated by the contraction of the Dirac matrices $\gamma^{\alpha}(-\gamma_{\alpha}\gamma_{\sigma}+\gamma_{\sigma}\gamma_{\alpha})/4=\gamma_{\sigma}\;(-\texttt{D}+(2-\texttt{D}))/4=\gamma_{\sigma}\;(-\texttt{D}+1)/2$. So the field virial of the spinors vanishes, hence the Dirac Lagrangian is conformal invariant in any dimension. By the same argument, the Rarita-Schwinger spin $3/2$ Lagrangian ${\overline{\psi}}_{\mu}\not{\partial}\psi_{\mu}$, where we used the constraint ${\sigma}\gamma^{\mu}\psi_{\mu}=0$ to eliminate the spin $1/2$ components, is also conformal invariant in any dimension, because the constraint eliminates the contribution of the vector index rotation.
We now examine the Maxwell Lagrangian
$\begin{array}[]{c}\mathcal{L}=-\frac{1}{4}(\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu})(\partial^{\mu}A^{\nu}-\partial^{\nu}A^{\mu})\;,\\ \\ V_{\sigma}(A)=\displaystyle{\frac{\partial\mathcal{L}}{\partial(\partial_{\alpha}A_{\mu})}}\;(\texttt{d}\;g_{\alpha\sigma}\;-\Sigma_{\alpha\sigma})\;(A_{\mu})\\ =-(\texttt{d}-1)A^{\alpha}(\partial_{\sigma}A_{\alpha}-\partial_{\alpha}A_{\sigma})\;.\end{array}$ (8)
The field virial vanishes only in dimension $(\texttt{d}=1\Leftrightarrow\texttt{D}=4)$, hence the Maxwell theory is conformal invariant only in 4 dimensions Bateman (1910); Jackiw and Pi (2011); El-Showk et al. (2011). Actually, it is also secretely invariant in $\texttt{D}=3$ because then the theory is degenerate and $F_{\mu\nu}$ is dual to a scalar (same references). In $\texttt{D}=4$ dimensions, we could turn in Yang-Mills interactions or include in the Lagrangian terms in $\Phi^{4}$, $A^{4}$, or $A^{2}\Phi^{2}$, this would introduce covariant derivatives on the right hand sides without altering the conformal invariance since, in the definition of the virial (3), we only differentiate relative to the derivatives of the fields. Yet mass terms are forbidden because they break the global scale invariance of the theory. In $\texttt{D}>4$ dimensions, the coupling constant g in the covariant derivative $D_{\mu}=\partial_{\mu}+\texttt{g}A_{\mu}$ has negative dimensions, also breaking conformal invariance.
3 Antisymmetric tensor fields
Antisymmetric tensor fields $T_{\mu\nu...}$ of rank p are equivalent to p-forms
$\begin{array}[]{c}T=(1/p!)\;T_{\mu\nu...}dx^{\mu}\wedge dx^{\nu}...\end{array}$ (9)
where $\wedge$ denotes the antisymmetric exterior product of the $1$-forms $dx^{\mu}$. They were first discussed in quantum field theory in relation with models for higher spin fields, see for example Ogievetsky and Polubarinov (1967), and they play an important role in supergravity and superstring theory. The most natural Lagrangian ($KR$), often associated in physics to the names of Kalb and Ramond (1974)
$\begin{array}[]{c}\mathcal{L}^{KR}\;d^{\texttt{D}}x=\frac{1}{2}({}^{*}dT)\wedge(dT)\end{array}$ (10)
is a straightforward generalization of the Maxwell Lagrangian $\mathcal{L}^{M}\;d^{4}x={\scriptstyle 1/2}\;{}^{*}dA\wedge dA$, where $A=A_{\mu}dx^{\mu}$ denotes the Maxwell 1-form, $d$ the Cartan exterior differential, the Hodge dual, and $d^{\texttt{D}}x$ the volume element D-form. Since $d$ is nilpotent, $d^{2}=0$, the Lagrangian $\mathcal{L}^{KR}$ is gauge invariant under a transformation of the p form $T$ parameterized by a $(\texttt{p}-1)$-form $\xi$:
$\begin{array}[]{c}\delta T^{(\texttt{p})}=d\xi^{(\texttt{p}-1)}\;\;\Rightarrow\;\;\delta\mathcal{L}^{KR}=0\;.\end{array}$ (11)
However, this is not the only possibility. In a stunning development, Avdeev and Chizhov ($AC$) discovered in 1994 a new Lagrangian describing in 4 dimensions a conformally invariant massless self-dual 2-tensor naturally interacting with chiral Fermions Avdeev and Chizhov (1994b, a). The phenomenology is well developed in Chizhov (2011).
Here, we study these 2 types of tensors in any dimension. We consider the Kalb-Ramond gauge tensor fields ($KR$) in dimension D, we generalize the known $\texttt{D}=4$ Avdeev-Chizhov matter tensor fields ($AC$) to higher dimensions, and we study the conformal invariance of these theories. In addition, we construct a new class of Lagrangians using a skew metric in the internal space that we call ($CP$) for two reasons. We prove that our new Lagrangian $\mathcal{L}^{CP}$ is $C$: conformal invariant and we show that it switches sign under the separate action of $C$: charge conjugation and $P$: parity conjugation, but remains invariant under the combined Landau $CP$: charge-parity conjugation.
The $\mathcal{L}^{CP}$ Lagrangian was first encountered as the tensor propagator in our recent study of the $SU(2/1)$ superalgebraic structure of the electroweak interactions Thierry-Mieg and Jarvis (2021). By checking in the next section the scalar-vector-tensor triple coupling present in that model, we will finally show that the complete $SU(2/1)$ super-chiral super-connection Lagrangian is conformal invariant in 4 dimensions.
Let us now consider the two standard tensor Lagrangians, Kalb-Ramond and Avdeev-Chizhov as provided as the very beginning of the second paper on this subject Avdeev and Chizhov (1994a) (and rediscovered in curved space by Shapiro (2024) equations 11 and 13)
$\begin{array}[]{c}\mathcal{L}^{KR}=\frac{1}{4}[(\partial_{\alpha}T_{\mu\nu})(\partial^{\alpha}T^{\mu\nu})-2(\partial^{\alpha}T_{\alpha\mu})(\partial_{\beta}T^{\beta\mu})]\;,\\ \mathcal{L}^{AC}=\frac{1}{4}[(\partial_{\alpha}T_{\mu\nu})(\partial^{\alpha}T^{\mu\nu})-4(\partial^{\alpha}T_{\alpha\mu})(\partial_{\beta}T^{\beta\mu})]\;.\end{array}$ (12)
Up to a total derivative, the trivial Lagrangian $T_{\mu\nu}\Box T^{\mu\nu}$ is among their linear combinations. But it is not conformal invariant because it treats the 6 components of $T_{\mu\nu}$ as six independent scalars without recognizing their tensorial spin structure. Therefore the Kalb-Ramond and the Avdeev-Chizhov Lagrangians cannot both be conformal invariant. Generalizing these equations, let $T$ be a p-form and let us compute the field virial associated to the d’Alembertian ($T\Box T)$ and to the square of the divergence (.$(\partial.T)^{2}$ in D dimension
$\begin{array}[]{c}\mathcal{L}^{\Box}=-\frac{1}{2}\;(T_{[\mu\nu...]})\;\Box(T^{[\mu\nu...]})\;,\\ V_{\sigma}^{\Box}=\texttt{d}(\partial_{\sigma}T_{[\mu\nu...]})\;T^{[\mu\nu...]}-2\texttt{p}\;(\partial_{\alpha}T^{[\alpha\nu...]})\;T_{[\sigma\nu...]}\;,\\ \\ \mathcal{L}^{\nabla}=\frac{1}{2}\;(\partial^{\alpha}T_{[\alpha\nu...]})\;(\partial_{\beta}T^{[\beta\nu...]})\;,\\ V_{\sigma}^{\nabla}=(\texttt{D}-\texttt{d}-\texttt{p})\;(\partial_{\alpha}T^{[\alpha\nu...]})\;T_{[\sigma\nu...]}\;.\end{array}$ (13)
The first term in $V_{\sigma}^{\Box}$ is a divergence and can be dropped, but the other 2 terms are not. However they can be combined. We can choose $k$ such that the linear combinations $V^{k}_{\sigma}=V_{\sigma}^{\Box}-\texttt{k}\;V_{\sigma}^{\nabla}$ of these two field virials becomes a divergence:
$\begin{array}[]{c}V^{\texttt{k}}_{\sigma}=V_{\sigma}^{\Box}-\texttt{k}\;V_{\sigma}^{\nabla}=\partial_{\sigma}(...)\;.\end{array}$ (14)
The necessary and sufficient condition is that that k solves
$\begin{array}[]{c}2\texttt{p}-\texttt{k}(\texttt{D}-\texttt{d}-\texttt{p})=0\;.\end{array}$ (15)
Therefore in this case and only in this case the corresponding Lagrangian $\mathcal{L}^{\texttt{k}}=\mathcal{L}^{\Box}-\texttt{k}\mathcal{L}^{\nabla}$ is conformally invariant. To apply this formalism, we work modulo integration by parts, denoted $\cong$, and rewrite the Maxwell-Kalb-Ramond Lagrangian in the desired form
$\begin{array}[]{c}\mathcal{L}^{KR}\;d^{\texttt{D}}x=\frac{1}{2}{}^{*}(dT)\wedge(dT)\cong\frac{1}{\texttt{p}!}(\mathcal{L}^{\Box}-\texttt{p}\mathcal{L}^{\nabla})\;d^{\texttt{D}}x\;.\end{array}$ (16)
[…]
Let us now analyze in more detail the Avdeev-Chizhov model. We know that in dimension $\texttt{D}=2p$, the antisymmetric p-tensor $T_{[\mu\nu...]}$ does not form an irreducible representation of the Lorentz group. Denoting ${}^{*}T$ the Hodge dual of the tensor $T$, we can split $T$ it into its Minkowski self-dual $Z$ and anti-self-dual part $\overline{Z}$ as follows
$\begin{array}[]{c}Z=T+i{}^{*}T\;,\;\;\overline{Z}=T-i{}^{*}T,\;\;{}^{*}Z=-iZ\;,\\ T=(Z+\overline{Z})/2\;,\;\;{}^{*}T=-i(Z-\overline{Z})/2\;.\end{array}$ (17)
These equations are valid in Minkowski space where ${}^{**}T=-T$. In Euclidean space, drop the $i$ and adjust the signs. Now, consider the chiral Lagrangian
$\begin{array}[]{c}\mathcal{L}^{\chi}=\frac{1}{2(p-1)!}(\partial.\overline{Z})(\partial.Z)=\frac{1}{2(\texttt{p}-1)!}(\partial^{\alpha}\overline{Z}_{\alpha\mu...})(\partial_{\beta}Z^{\beta\mu...})\;.\end{array}$ (18)
Using integration by parts, we can again rewrite this Lagrangian in the desired form
$\begin{array}[]{c}\mathcal{L}^{\chi}\cong\frac{1}{\texttt{p}!}(\mathcal{L}^{\Box}-\texttt{D}\mathcal{L}^{\nabla})\end{array}$ (19)
which coincides in $\texttt{D}=4$ dimensions with the Avdeev-Chizhov Lagrangian (12) and provides a generalization in any even dimension. So $\texttt{k}=\texttt{D}=2\texttt{p}$ and $\texttt{d}=(\texttt{D}-2)/2=\texttt{p}-1$, hence condition (15) $2\texttt{p}-\texttt{k}(\texttt{D}-\texttt{d}-\texttt{p})=0$ holds and therefore the generalized Avdeev-Chizhov theory is conformal invariant in any even dimension. We recover $\texttt{p}=1$ in $\texttt{D}=2$, i.e. the pseudo-scalars of string theory, $\texttt{p}=2$ in $\texttt{D}=4$, i.e. the original Avdeev-Chizhov Lagrangian, and so on.
This simple structure of the chiral Lagrangian $\mathcal{L}^{\chi}$ provides a way to construct the antisymmetric pseudo-scalar Lagrangian whose existence was proven in Jarvis and Thierry-Mieg (2023) using the analysis of the Lorentz group. Let us add an internal index to the tensors. If the internal metric $\kappa_{ab}$ of the internal space is symmetric, we recover the generalized charged Avdeev-Chizhov Lagrangian
$\begin{array}[]{c}\mathcal{L}^{AC}=\frac{1}{2(p-1)!}\;\kappa_{ab}(\partial.\overline{Z}^{a})(\partial.Z^{b})\\ =\frac{1}{\texttt{p}!}\;\kappa_{ab}(-T^{a}\Box T^{b}-\texttt{D}(\partial.T^{a})(\partial.T^{b}))\;,\;\;\kappa_{ab}=\kappa_{ba}\;.\end{array}$ (20)
But we can also construct an antisymmetric Lagrangian, Thierry-Mieg and Jarvis (2021), using an antisymmetric metric $\eta_{ij}$ in the internal space
$\begin{array}[]{c}\mathcal{L}^{CP}=\frac{-i}{(p-1)!}\;\eta_{ij}(\partial.\overline{Z}^{i})(\partial.Z^{j})\\ =\;\frac{1}{(p-1)!}\;\eta_{ij}(\partial.{}^{*}T^{i})(\partial.T^{j})\;,\;\;\eta_{ij}=-\eta_{ji}\;.\end{array}$ (21)
This Lagrangian is real, because complex conjugation switches the sign of the overall factor $i$ but at the same time exchanges $Z$ and $\overline{Z}$ in the definition (17). It is pseudo-scalar, because flipping the orientation of space exchanges $(Z,\overline{Z})$ and therefore flips the sign of the Lagrangian. Yet it naturally implements the Landau $CP$ symmetry because, if the orientation of the internal space, the indices $[ij]$, are flipped at the same time as the orientation of space, the sign of the Lagrangian is maintained. We call it $CP$ because it implements Landau charge-parity $CP$ symmetry, and also because it is both $C$:conformal invariant (see below) and $P$:parity aware.
In its $Z$ form, the left hand side of the $\mathcal{L}^{AC}$ (20) is very similar to $\mathcal{L}^{CP}$ (21), but when we rewrite it as a function of $T$, it is markedly different. In particular the d’Alembertian term, which is necessarily symmetric, disappears. Thus, to prove the conformal invariance of our Lagrangian $\mathcal{L}^{CP}$, we cannot use the field virials $V^{\Box}$ and $V^{\nabla}$ defined in (13). To compute the $V(Z)$ field virial, we start again from the definition (3). The advantage is that this new calculation will cover both the symmetric $\mathcal{L}^{AC}=\mathcal{L}^{\chi}_{\kappa}$ and the antisymmetric $\mathcal{L}^{CP}=\mathcal{L}^{\chi}_{\eta}$ cases. Using $\partial\overline{Z}/\partial Z=0$, a direct a consequence of self-duality reminiscent of complex derivatives, we find
$\begin{array}[]{c}V^{\chi,\nabla}_{\sigma}=\displaystyle{\frac{\partial\mathcal{L}^{\chi}}{\partial(\partial_{\alpha}\overline{Z}^{i}_{\beta\gamma...})}}\;(\delta g_{\alpha\sigma}-\Sigma_{\alpha\sigma})(\overline{Z}^{i}_{\beta\gamma...})+\\ +\displaystyle{\frac{\partial\mathcal{L}^{\chi}}{\partial(\partial_{\alpha}Z^{j}_{\beta\gamma...})}}\;(\delta g_{\alpha\sigma}-\Sigma_{\alpha\sigma})(Z^{j}_{\beta\gamma...})\\ =\frac{1}{\texttt{p}!}(\texttt{d}-\texttt{D}+\texttt{p})\;\eta_{ij}\;[(\partial_{\alpha}\overline{Z}^{i,\alpha\mu...})\;Z^{j}_{\sigma\mu...}+(\partial_{\alpha}Z^{j,\alpha\mu...})\;\overline{Z}^{i}_{\sigma\mu...}]\;.\end{array}$ (22)
We wish to establish that this field virial is a divergence. To complete the proof, we take advantage of an interesting identity. Consider the d’Alembertian $\Box$ of the chiral fields
$\begin{array}[]{c}\mathcal{L}^{\chi,\Box}=-\frac{1}{2\texttt{p}!}\;\eta_{ij}\;\overline{Z}^{i,\mu\nu...}\;\Box\;Z^{j}_{\mu\nu...}\end{array}$ (23)
In reality, this Lagrangian vanishes because the total contraction of an anti self-dual $\overline{Z}$ with a self-dual tensor $Z$ vanishes. Now since $Z$ is self-dual, and since $\Box$ does not alter the tensorial indices, $\Box Z$ is also self-dual and $\mathcal{L}^{\chi,\Box}=0$. Nevertheless, we can compute the corresponding field virial. Obviously, it vanishes up to a total derivative, but in a very non trivial way.
$\begin{array}[]{c}\par V^{Z,\Box}_{\sigma}=-\;\frac{1}{\texttt{p}!}\;\eta_{ij}\;(\overline{Z}^{i,\mu\nu...})\;\partial^{\alpha}\;(\texttt{d}g_{\alpha\sigma}-\Sigma_{\alpha\sigma})(Z^{j}_{\mu\nu...})\\ =-\frac{1}{\texttt{p}!}\;\eta_{ij}(\overline{Z}^{i,\mu\nu...})\;(\texttt{d}\partial_{\sigma}Z^{j}_{\mu\nu...}+\texttt{p}(-\partial_{\mu}Z^{j}_{\sigma\nu...}+g_{\mu\sigma}\partial^{\alpha}Z_{\alpha\nu...}))\;.\end{array}$ (24)
The first term ($\eta\;\overline{Z}\;\texttt{d}\;\partial_{\sigma}Z$) vanishes (self-dual to anti-self-dual full contraction), but the second term does not, so
$\begin{array}[]{c}V^{Z.\Box}_{\sigma}\cong\frac{1}{(\texttt{p}-1)!}\;\eta_{ij}(\overline{Z}^{i,\mu\nu...}\;\partial_{\mu}\;Z^{j}_{\sigma\nu...}-\overline{Z}^{i}_{\sigma\nu...}\;\partial_{\mu}Z^{j,\mu\nu...})\;.\end{array}$ (25)
Adding the $\overline{Z}$ derivative we get
$\begin{array}[]{c}V^{\chi,\Box}_{\sigma}=V^{\overline{Z},\Box}_{\sigma}+V^{Z,\Box}_{\sigma}\\ =\frac{1}{(\texttt{p}-1)!}\;\eta_{ij}\;(\overline{Z}^{i,\mu...}\;\partial_{\mu}\;Z^{j}_{\sigma...}-\overline{Z}^{i}_{\sigma...}\;\partial_{\mu}Z^{j,\mu...}+\\ +Z^{j,\mu...}\;\partial_{\mu}\;\overline{Z}^{i}_{\sigma...}-Z^{j}_{\sigma...}\;\partial_{\mu}\overline{Z}^{i,\mu...})\;.\end{array}$ (26)
Adding the 2 types of field virials with the proper linear combination, we get a divergence
$\begin{array}[]{c}(\texttt{d}-\texttt{D}+\texttt{p})V^{\chi,\Box}_{\sigma}+2\texttt{p}\;V^{\chi,\nabla}_{\sigma}=\\ (\texttt{d}-\texttt{D}+\texttt{p})\;\frac{1}{(\texttt{p}-1)!}\;\eta_{ij}\;\partial_{\mu}(\overline{Z}^{i}_{\sigma...}Z^{j,\mu...}+\overline{Z}^{i,\mu...}Z^{j}_{\sigma...})\;.\end{array}$ (27)
Since we know a priori that the $V^{\chi,\Box}$ field virial vanishes up to a divergence for the simple reason that the Lagrangian $\mathcal{L}^{\chi,\Box}$ vanishes identically, we conclude that $V^{\chi,\nabla}_{\sigma}$ is a divergence, and therefore the Lagrangian $\mathcal{L}^{\chi,\nabla}$ is conformal invariant for any choice of the internal metric. In the symmetric case $\mathcal{L}^{AC}=\mathcal{L}^{\chi,\nabla}_{\kappa}$ where $\kappa_{ab}=\kappa_{ba}$, we recover the already demonstrated conformal invariance of the Avdeev-Chizhov Lagrangian. In the antisymmetric case $\mathcal{L}^{CP}=\mathcal{L}^{\chi,\nabla}_{\eta}$, where $\eta_{ij}=-\eta_{ji}$, we prove that the $CP$ model is conformal invariant.
Fulfilling the classification Jarvis and Thierry-Mieg (2023), there exists a second pseudo-scalar Lagrangian, that we can call pseudo-trivial ($PT$)
$\begin{array}[]{c}\mathcal{L}^{PT}\;d^{\texttt{D}}x=-\frac{1}{2}\;\kappa_{ab}\;T^{a}\;\Box\wedge T^{b}\\ =-\frac{1}{2p!}\;\kappa_{ab}\;\epsilon^{\mu\nu...\rho\sigma...}\;T^{a}_{\mu\nu...}\;\Box\;T^{b}_{\rho\sigma...}\;d^{\texttt{D}}x\;.\end{array}$ (28)
This Lagrangian treats the $T_{\mu\nu...}$ component fields as a collection of scalars. Like $T^{\mu\nu...}\Box T_{\mu\nu...}$, it does not recognize the spin of the tensor $T$ and for this reason is not conformal invariant. In $\texttt{D}=4k$, it is symmetric in $(\mu\nu...,\rho\sigma...)$, so $\kappa_{ab}$ has to be symmetric in $(ab)$, skew in $\texttt{D}=4k+2$.
To summarize, we have constructed in D dimensions 4 different families of Lagrangians. Two are invariant ($KR$ and $AC$) and two flip sign if we change the orientation of space ($CP$ and $PT$). Two are using a symmetric metric $\kappa_{\{ab\}}$ in the internal space ($KR$, $AC$) and one an antisymmetric $\eta_{[ij]}$ metric ($CP$), finally $PT$ alternates. Two are conformal invariant ($AC$ and $CP$) in any even dimension and one ($KR$) if and only if the rank p of the tensor matches its canonical dimension d. These four cases saturate a Lorentz group-theoretical classification in $\texttt{D}=4$ dimensions (see Jarvis and Thierry-Mieg (2023), appendix A) which proved that there are no other possibilities. The corresponding character analysis using representations of the full Lorentz group in arbitrary even dimension D = 2p shows that there remain four one-dimensional candidates, namely two $O(2\texttt{p}\!-\!1,1)$ scalars in symmetric coupling, plus one $O(2\texttt{p}-1,1)$ pseudoscalar in each of symmetric and antisymmetric coupling. So the classification is complete in all even dimensions.
4 Physical content of the anti-symmetric tensor theories
[…]
$\begin{array}[]{c}\sum_{i=0}^{\texttt{p}}\;(-1)^{i}(i+1)(^{\texttt{p}-i}_{\texttt{D}})=(^{\texttt{p}}_{\texttt{D}-2})\;.\end{array}$ (29)
which coincides with the development of an anti-symmetric Young tableau of the BRS extended orthosymplectic superalgebra $OSp(D/2)$ of Delbourgo and Jarvis (1983) relative to its maximal even subalgebra $SO(D)\oplus Sp(2)$.
The situation is more complicated in the Avdeev-Chizhov case (20, 21). The tensors $Z$ and $\overline{Z}$ are self dual and anti-self-dual (17) and each have D!/ 2(($\texttt{D}!\;/\;2((\texttt{D}/2)!)^{2}$ components, i.e. 3 components in 4 dimensions. The free Lagrangian is invertible and the propagators are well defined. This is consistent with the absence of gauge invariance, since there is no need and no possibility of adding a gauge fixing term like in the Maxwell or Kalb-Ramond models. Nevertheless, in 4 dimensions, a real Avdeev-Chizhov tensor carries only one degree of freedom (Avdeev and Chizhov (1994a); Lemes et al. (1995a, b); Wetterich (2008))
Although there is no gauge invariance, Avdeev and Chizhov (1994a) proposed in 4 dimensions a possible set of Faddeev-Popov ghosts and BRS equations used by Kirilova and Chizhov (2017) in a way reminiscent of Sezgin and Van Nieuwenhuizen (1980). An alternative could be to consider again the BRS extended superalgebra and introduce the infinite tower of ghosts associated to the $OSp(4/2)$ self-dual supertensor consisting $\forall n\in\mathbb{N}$ of a self-dual 2-tensor at level $4n$, an anti-self-dual 2-tensor at level $24n+$ 2 and a ghost vector at level $12n+$ 1 with regularized sum $3-4+(3+1)-4+(3+1)-4\;...=3-4/2=1$. More generally the superdimension of the $OSp(D/2)$ self-dual supertensor is equal to the dimension of the self-dual tensor of the space $SO(D-2)$ transverse to the light cone (Stoilova et al. (2017)). This is most likely the number of physical components of the self dual model in dimension $D$, although an explicit Hamiltonian proof is lacking in $\texttt{D}>4$.
The phenomenology of self-dual tensors in $\texttt{D}=4$ dimensions is well detailed in Chizhov (2011). Their renormalization is analyzed in Lemes et al. (1995a, b) and Wetterich (2008). As shown in the interesting paper of Shapiro (2024), the curved-space extension of the Avdeev-Chizhov Lagrangian is locally scale invariant (Weyl conformal invariant). See also the geometric constructions of Hughes and Kusmartsev (2024).
The $CP$ model might be a promising candidate to construct an extension of the standard model because in 4 dimensions, self-dual 2-tensors naturally couple to chiral Fermions. Indeed the minimal coupling of an anti-symmetric 2-tensor $T_{[\mu\nu]}$ to a massless chiral spinor in $4$ dimensions $\overline{(\psi_{R})}_{L}T_{[\mu\nu]}{\sigma}^{\mu}{\overline{\sigma}}^{\nu}\psi_{L}$ automatically selects a self dual tensor because the antisymmetric product of 2 Pauli matrices is self dual ${}^{*}({\sigma}{\overline{\sigma}})={\sigma}{\overline{\sigma}}$. Furthermore, as shown in Thierry-Mieg and Jarvis (2021) the counter-term to the propagator of the $T$ tensor induced by the Fermion loops is constrained by the Bouchiat-Iliopoulos-Meyer (BIM) mechanism to induce an antisymmetric metric $\eta_{ij}$ in the internal space, thus favoring the (CP) model over the (AC) model. The Fermion loops also induce a triple scalar-vector-tensor vertex $\Phi\;F^{\mu\nu}T_{\mu\nu}$. As this term depends on the derivatives of the vector field $F_{\mu\nu}=\partial_{[\mu}A_{\nu]}$, it could contribute to the field virial and break conformal invariance. However
$\begin{array}[]{c}\mathcal{L}^{\Phi AB}=\frac{1}{2}\Phi F^{\mu\nu}T_{\mu\nu}\;,\\ V_{\sigma}^{CP}=\displaystyle{\frac{\partial\mathcal{L}^{\Phi AT}}{\partial(\partial_{\alpha}A_{\mu})}}\;(\delta\;g_{\alpha\sigma}\;-\Sigma_{\alpha\sigma})\;(A_{\mu})=(\texttt{d}-1)\Phi T_{\sigma\mu}A^{\mu}\;.\end{array}$ (30)
Hence the scalar-vector-tensor vertex $\Phi\;F^{\mu\nu}T_{\mu\nu}$ is conformally invariant in $\texttt{d}=1\Leftrightarrow\texttt{D}=4$ dimensions.
5 Conclusion
Conformal invariance is an important symmetry (see for example the physics report of Nakayama (2015)), and we provide for the first time a systematic analysis of the conformal invariance of antisymmetric-tensors Lagrangian field theories in any even dimension. First, we have analyzed the Kalb-Ramond gauge tensor Lagrangian and using the conformal field virial $V_{\sigma}$ (3) constructed by Jackiw and Pi (2011), we have shown (16) that the theory is conformal invariant if and only if the rank p of the gauge tensor is equal to its canonical dimension $\texttt{d}=(\texttt{D}-2)/2$. This case covers the $\texttt{p}=0$ scalars of the 2-dimensional string theory, the $\texttt{p}=1$ vector of the 4-dimensional Maxwell-Yang-Mills theory, the $\texttt{p}=2$ Kalb-Ramond 2-tensor theory in 6 dimensions with cubic interactions, and also 3-tensors in 8 dimensions, 4-tensors in 10 dimensions and so on. This direct proof is new, compare with Kuzenko and Turner (2021); Lee and Zheng (2021). However, the Kalb-Ramond 2-form gauge fields are not conformal invariant in 4 dimensions, and the Maxwell theory is conformal invariant only in 4 dimensions, Bateman (1910).
We then analyzed the Avdeev and Chizhov (1994b) matter tensor Lagrangian (12), generalized it to higher dimensions and proved for the first time that the self-dual matter tensor fields are conformal invariant in any even dimension (19).
Furthermore, by casting the theory into a chiral form (20), we constructed in any even dimension a novel family of Lagrangians (21) describing charged self-dual tensors which we call $\mathcal{L}^{CP}$ because they implement the Landau charge-parity $CP$ invariance, and because we proved that these Lagrangians are $C$:conformal invariant (27) and $P$:parity aware. We first discovered these new tensors in our analysis of the superalgebraic $SU(2/1)$ structure of quarks and leptons. By analyzing the novel scalar-vector-tensor triangle interaction induced in that theory by the Fermion loop counter-terms, we prove (30) that the complete $SU(2/1)$ super-chiral super-connection Lagrangian we constructed in Thierry-Mieg and Jarvis (2021) is conformally invariant, reinforcing the possibility that it could play a role in a superalgebraic extension of the standard model.
We hope that this focused study on conformal invariance, based on the elegant construction of Jackiw and Pi, will raise interest in self-dual tensors.
Acknowledgments
We are gratefully to Sergei Kuzenko, Ivan Todorov, and especially Mikhail Chizhov, for comments. We also thank the referees for their useful suggestions. This research was supported in part by the Intramural Research Program of the National Library of Medicine, National Institute of Health.
.1 The gauged fixed Maxwell model is conformal invariant in $\texttt{D}\geq 4$
It is well known that the Maxwell theory is conformally invariant in $\texttt{D}=4$ dimensions, Bateman (1910), or in $\texttt{D}=3$ by duality, but not in $\texttt{D}>4$ (see for example Jackiw and Pi (2011); El-Showk et al. (2011)). However, in the same paper Showk noticed noticed that the gauged fixed Maxwell model is conformal invariant in $\texttt{D}>4$ for a particular gauge choice. That proof can be greatly simplified using the Jackiw and Pi theorem as follows. Consider the gauged fixed Maxwell Lagrangian
$\begin{array}[]{c}\mathcal{L}^{MGF}=\mathcal{L}^{M}+\mathcal{L}^{GF}\;,\;\;\mathcal{L}^{M}=-\frac{1}{4}(F_{\mu\nu})^{2}\;,\;\;\mathcal{L}^{GF}=-\frac{1}{2\xi}\;(\partial^{\mu}A_{\mu})^{2}\;.\end{array}$ (31)
We dropped the Faddeev-Popov ghosts, since scalars are conformal invariant. The Jackiw current associated to the classical Maxwell term has already by computed (8)
$\begin{array}[]{c}V^{M}_{\sigma}(A)=\frac{\partial\mathcal{L}^{M}}{\partial(\partial_{\alpha}A_{\mu})}\;(\texttt{d}\;g_{\alpha\sigma}\;-\Sigma_{\alpha\sigma})\;(A_{\mu})=-(\texttt{d}-1)A^{\alpha}(\partial_{\sigma}A_{\alpha}-\partial_{\alpha}A_{\sigma})\;.\end{array}$ (32)
The first term is by itself a divergence $A^{\alpha}\partial_{\sigma}A_{\alpha}=1/2\;\partial^{\beta}(\delta_{\beta\sigma}\;A^{\alpha}A_{\alpha})$, but the second term, which can be rearranged as $(\texttt{d}-1)(\partial^{\alpha}A_{\alpha})A_{\sigma}$ is not. However, we must also consider the Jackiw current $V^{GF}$ associated to the gauge fixing term
$\begin{array}[]{c}V^{GF}_{\sigma}=\frac{\partial\mathcal{L}^{GF}}{\partial(\partial_{\alpha}A_{\mu})}\;(\texttt{d}\;g_{\alpha\sigma}\;-\Sigma_{\alpha\sigma})\;(A_{\mu})=-\frac{1}{\xi}(\texttt{d}+1-\texttt{D})(\partial^{\mu}A_{\mu})\;A_{\sigma}\end{array}$ (33)
Therefore if $\xi=(\texttt{D}-\texttt{d}-1)/(\texttt{d}-1)=\texttt{D}/(\texttt{D}-4)$, the unwanted second term is canceled, the total Jackiw current $J^{M}+J^{GF}$ is a divergence, and the model is conformal invariant. We have recovered the curious observation of Showk, but it a much simpler way. Notice however that conformal invariance only applies to the free theory. The canonical dimension of the vector field $A$ is $\texttt{d}=(\texttt{D}-2)/2$ so $d=1$ only in dimension $D=4$. Therefore the coupling constant g in the covariant derivative $D_{\mu}=\partial_{\mu}+\texttt{g}\;A_{\mu}$ is dimensionless only in $\texttt{D}=4$, hence the gauge fixed coupled QED or Yang-Mills models are not conformal invariant in $\texttt{D}>4$.
.2 The quartic D=6 vector model is not conformal invariant
In section C.3 of Lee and Zheng (2021), it is claimed that in 6 dimensions, the quartic Lagrangian
$\begin{array}[]{c}\mathcal{L}^{Q}=-\frac{1}{4}\;F^{\mu\nu}\Box F_{\mu\nu}\;,\;\;\;F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}\;,\end{array}$ (34)
is conformal invariant. The advantage of this quartic Lagrangian is that in $\texttt{D}=6$ the $A_{\mu}$ vector field has canonical dimension 1, allowing covariant derivatives with a dimensionless coupling constant, but using Jackiw and Pi we can show that this Lagrangian is not conformal invariant because the $\Box$ d’Alembertian does not take into account the tensorial nature of $F_{\mu\nu}$ and treats it as a collection of 6 scalars. Indeed, when we compute the derivative $\partial/\partial(\partial_{\alpha}A_{\mu})$, the $\Box$ term is not inert but contributes to the current and we find
$\begin{array}[]{c}V^{Q}_{\sigma}=-(\texttt{d}-1)(\partial_{\sigma}\partial_{\mu}(\partial.A))\;A_{\mu}+\texttt{d}\;(\partial_{\sigma}\;\Box A_{\mu})\;A_{\mu}-(\partial_{\mu}\;\Box A_{\sigma})\;A_{\mu}\end{array}$ (35)
which is not a divergence. Hence the quartic model $\mathcal{L}^{Q}$ is not conformal invariant.